# A Brain-to-Population Graph Learning Framework for Diagnosing Brain Disorders

### v1

@@ Abstract
Figure 1: The framework of our proposed Brain-to-Population Graph Learning Network. The upper section depicts the first stage, which constructs semantic-enhanced brain graphs and learns brain representations using Adaptive Node Reassignment Graph Attention Network (ANR-GAT). The lower section illustrates the second stage, building a condition-based population graph and capturing meaningful patterns via a Heterogeneous Graph Convolutional Network (HGCN) and a gated fusion mechanism to integrate phenotypic data.
@@ 1 Introduction
Numerous computer-aided diagnosis methods have been developed based on FC analysis for diagnosing neurological disorders. Among these, graph neural networks (GNNs) have become a dominant approach owing to their powerful graph embedding capabilities. Building on this strength, recent works (Scarselli et al., 2008) and (Cui et al., 2022) have demonstrated the effectiveness of GNNs in analyzing both brain networks and population networks. However, two main limitations exist in current brain network GNNs: (1) the underutilization of essential information encoded in brain atlases, and (2) the sole focus on individual FC embeddings, neglecting the inter-individual relationships both within and across groups (Wang et al., 2024b). Although the population GNNs incorporate non-imaging data to model relationships between subjects, they often rely on linking individuals from the same sites or with similar demographics (Huang and Chung, 2020; Zhang et al., 2023a). Given that GNNs operate through neighborhood aggregation (Chen et al., 2024a), this approach amplifies the biases inherent in FC data, where confounding factors such as site and demographic variations already mask disease-relevant patterns.
@@ 2.1 GNN-based Neurological Disorders Diagnosis
Researchers have explored multi-view FC across different atlases (Wang et al., 2024b), multi-modal MRI data (Chen et al., 2022b; Zhu et al., 2022), and dynamic FC (Campbell et al., 2023) to enrich FC representations in brain networks. However, these methods often require additional modalities or costly preprocessing. To improve brain network representation, advanced graph structures like hypergraphs (Ji et al., 2022; Zuo et al., 2024) and heterogeneous graphs (Wen et al., 2024), as well as learning methods such as metric learning (Ktena et al., 2018), contrastive learning (Zhang et al., 2023b; Wang et al., 2025), and adversarial learning (Chen et al., 2022b; Zuo et al., 2024), have been applied to capture key patterns in FC data. In our brain representation learning, we enhance brain network representation by leveraging large language models (LLMs) to extract rich textual descriptions of brain regions, encoding their functional and structural characteristics. These embeddings provide additional semantic information, enriching the graph representations in brain network analysis.
In population networks, neurological disorder prediction is treated as a node classification task. Node features are derived from the upper triangle of FC matrices (Shao et al., 2023) or extracted using methods like recursive feature elimination (RFE) (Yang et al., 2023). Alternatively, GNNs are employed to learn graph representations of FC for node feature embedding (Jiang et al., 2020; Zhang et al., 2023a). Non-imaging data, such as phenotypic features, are integrated by mapping them to edge weights (Huang and Chung, 2020; Zhang et al., 2023a) or learning modality-shared and modality-specific representations to construct population networks (Zheng et al., 2022; Chen et al., 2024b). While some recent studies have explored jointly modeling brain-level and population-level graphs to capture both within- and between-subject patterns (Pan et al., 2022; Li and Zhang, 2024; Zeng et al., 2025), we use a two-stage approach to help disentangle subject-specific neural features from broader population-level variations.
@@ 3.1 Weighted Graph Attention in Brain Network
[…]
@@ 3.2 Heterogeneous Graph Convolution in Population Network
The population graph is defined as $G_{p}=(\mathcal{V}_{p},\mathcal{E}_{p},\bm{X}_{p},\mathcal{C})$, where $\mathcal{V}_{p}$ represents individual nodes, $\mathcal{E}_{p}$ represents population edges, $\bm{X}_{p}\in\mathbb{R}^{|\mathcal{V}_{p}|\times F_{2}}$ is the node feature matrix with $F_{2}$ denoting the feature dimension and $\mathcal{C}$ is a set of conditions that define the relationships between nodes. A heterogeneous graph network (HGCN) (Schlichtkrull et al., 2018; Zhang et al., 2019) is used to propagate information across different relations. Each edge corresponds to a unique relationship under a specific condition $c\in\mathcal{C}$.
@@ 4.1.1 Semantic-Enhanced Brain Graph
In this work, we leverage large language models (LLMs) to derive semantic embeddings of brain regions, capturing both their functional roles and structural properties. For each brain region, prompts are constructed to elicit descriptive text capturing its core neurobiological attributes. These responses are then tokenized and embedded using the text encoder. An illustration of this process is provided in Figure 2. To quantify the semantic relationships between regions, we compute pairwise cosine similarity between their embeddings: $\text{Sim}(u,v)=\frac{\mathbf{s}_{u}\cdot\mathbf{s}_{v}}{|\mathbf{s}_{u}||\mathbf{s}_{v}|}$, where $\mathbf{s}_{u}$ and $\mathbf{s}_{v}$, where $\mathbf{s}_{u}$ and $\mathbf{s}_{v}$ denote the semantic embeddings of brain regions $u$ and $v$, respectively. Based on these similarity scores, we construct edges in the brain graph, using a thresholded similarity score as the edge weight. This incorporation of semantic similarity enables the graph structure to more faithfully reflect the functional and structural relevance between regions.
@@
To ensure consistency and reproducibility in concept generation, we designed a structured prompt format rooted in established neuroscience standards, such as the Terminologia Neuroanatomica and the Human Connectome Project. The schema guiding the initial neuroanatomical descriptions is illustrated in Figure 3.
@@
To further enhance the granularity and generalizability of the extracted features, we introduce a secondary prompt tailored to extract hierarchical, ontology-driven concepts for each brain region. This prompt guides the model to produce fine-grained annotations that align with known structural-functional hierarchies. An overview of this concept extraction process is shown in Figure 4.
@@ 4.1.2 Feature-to-Group Mapping
As part of the ANR-GAT model, the node feature matrix $\bm{H}_{b}\in\mathbb{R}^{|\mathcal{V}_{b}|\times F^{\prime}}$ is obtained after the WGAT layer. To reassign nodes into new groups based on their feature representations, we introduce a feature-to-group mapping mechanism that projects these features into a shared group space. The group assignment matrix $\bm{M}\in\mathbb{R}^{|\mathcal{V}_{b}|\times N_{g}}$, where $N_{g}$ is the number of groups, is computed by applying a learnable reassignment weight matrix $\bm{W}_{r}\in\mathbb{R}^{F^{\prime}\times N_{g}}$ followed by a softmax operation: $\bm{M}=\text{Softmax}(\bm{H}_{b}\bm{W})$, where the softmax is applied row-wise to ensure that the group assignment probabilities for each node sum to one.
For each node $i\in\mathcal{V}$, the group assignment $\text{group}_{i}$ is determined by selecting the group with the highest probability from $\bm{M}$ along the group dimension: $\text{group}_{i}=\arg\max_{j}M_{ij},\quad\forall j\in\{1,2,\dots,N_{g}\}.$ As illustrated in Figure 1, this mapping is a key component of the ANR-GAT model, designed for brain graph representation learning. Node features are aggregated within each group, combining information from nodes reassigned to the same group. The edges are also remapped between groups, guided by the original graph connectivity and the reassigned groups. This process simplifies the graph structure while preserving key relationships, enabling the generation of new node representations that are passed to subsequent layers of WGAT for further processing. As an intermediate step bridging the WGAT layers, the feature-to-group mapping facilitates hierarchical feature learning and graph structure simplification.
@@ 4.3 Two-Stage Optimization for B2P-GL Framework
The proposed framework is comprised of two stages. The first stage is devoted to optimizing the representation of the brain graph, while the second stage aims to enhance the classification of disorders.
@@ 5.1 Datasets
Table 1 summarizes the demographic characteristics of the three datasets. ADHD-200 primarily includes children and adolescents, whereas ABIDE I and REST-meta-MDD cover a broader age range extending into adulthood. ABIDE I and ADHD-200 show a clear male predominance, while REST-meta-MDD has a more balanced gender distribution. All datasets are collected from multiple sites, contributing to demographic and acquisition variability. ABIDE I includes 850 participants from 20 sites, ADHD-200 consists of 938 participants from 7 sites, and the REST-meta-MDD comprises 1,567 participants across 16 sites. Both ABIDE I and ADHD-200 provide Full-Scale, Performance, and Verbal IQ scores, although a substantial portion of these values are missing in ABIDE I and ADHD-200 datasets. REST-meta-MDD includes years of education, which may serve as a proxy for cognitive and socioeconomic background.
@@ 5.2 Experimental Settings
Models were implemented and evaluated within a PyTorch environment using an A800 GPU. Table 2 summarizes the hyperparameter configurations for the two stages of the B2P-GL model across the ABIDE I, ADHD-200, and REST-meta-MDD datasets. The first stage adopts the ANR-GAT architecture, while the second stage is based on the HGCN framework. The table outlines dataset-specific settings for each stage, including the optimizer, learning rate, batch size, number of training epochs, hidden dimensions, and selected phenotypic variables.
@@
learning rate $1.00\times 10^{-4}$ $1.00\times 10^{-4}$ $1.00\times 10^{-4}$
@@
learning rate $1.00\times 10^{-2}$ $1.00\times 10^{-2}$ $1.00\times 10^{-2}$
@@
5.3 Quantitative Comparison of Model Performance
@@ 5.4.1 Effectiveness of Semantic Similarity in Edge
In addition, we further evaluated two strategies for computing semantic similarity: (1) lexical embeddings of region names and (2) contextualized embeddings incorporating their concepts. Subsequently, we compared BERT-base (Devlin et al., 2018), a pre-trained language model designed for general tasks, and BioBERT, a biomedical domain-adapted variant. As shown in Table 5, incorporating contextualized concepts led to consistent performance gains. For BERT-base, concept-based embeddings improved accuracy by 0.6%. BioBERT also benefited from concept augmentation, with a 0.4% increase. When comparing models under the same input setting (i.e., concepts), BioBERT outperformed BERT-base by 0.3% in accuracy and 0.4% in AUC, indicating its advantage as a more suitable choice for neurological applications. These results highlight the importance of both domain-adapted language models and concept-level augmentation for capturing meaningful semantic similarity between brain regions.
@@
Names BERT 72.8 ± 0.6 72.3 ± 0.6 80.2 ± 1.8 64.4 ± 2.4
BioBERT 73.3 ± 0.8 72.9 ± 0.8 79.2 ± 1.5 66.5 ± 1.6
Concepts BERT 73.4 ± 0.5 72.8 ± 0.6 81.7 ± 2.0 64.0 ± 2.5
BioBERT 73.7 ± 0.6 73.2 ± 0.6 80.5 ± 1.8 65.9 ± 2.1

### v2

@@ Abstract
Figure 1: The framework of our proposed Brain-to-Population Graph Learning Network. The upper section depicts the first stage, which constructs semantic-enhanced brain graphs and learns brain representations using the Adaptive Node Reassignment Graph Attention Network (ANR-GAT). The lower section illustrates the second stage, building a condition-based population graph and capturing meaningful patterns via a Heterogeneous Graph Convolutional Network (HGCN) and a gated fusion mechanism to integrate phenotypic data.
@@ 1 Introduction
Numerous computer-aided diagnosis methods have been developed based on FC analysis for diagnosing neurological disorders. Among these, graph neural networks (GNNs) have become a dominant approach owing to their powerful graph embedding capabilities. Building on this strength, recent works (Scarselli et al., 2008) and (Cui et al., 2022) have demonstrated the effectiveness of GNNs in analyzing both brain networks and population networks. However, two main limitations exist in current brain network GNNs: (1) the underutilization of essential information encoded in brain atlases, and (2) the sole focus on individual FC embeddings, neglecting the inter-individual relationships both within and across groups (Wang et al., 2024b). Although the population GNNs incorporate non-imaging data to model relationships between subjects, they often rely on linking individuals from the same sites or with similar demographics (Huang and Chung, 2020; Zhang et al., 2023a). Given that GNNs operate through neighborhood aggregation (Chen et al., 2024a), this approach amplifies the biases inherent in FC data, where confounding factors such as site and demographic variations already mask disease-relevant patterns and may make it harder for the model to learn class-discriminative features (Kazi et al., 2019).
@@ 2.1 GNN-based Neurological Disorders Diagnosis
Researchers have explored multi-view FC across different atlases (Wang et al., 2024b), multi-modal MRI data (Chen et al., 2022b; Zhu et al., 2022), and dynamic FC (Campbell et al., 2023) to enrich FC representations in brain networks. However, these methods often require additional modalities or costly preprocessing. To improve brain network representation, advanced graph structures like hypergraphs (Ji et al., 2022; Zuo et al., 2024) and heterogeneous graphs (Wen et al., 2024), as well as learning methods such as metric learning (Ktena et al., 2018), contrastive learning (Zhang et al., 2023b; Shi et al., 2024; Wang et al., 2025), self-supervised learning (Shi et al., 2025), and adversarial learning (Chen et al., 2022b; Zuo et al., 2024), have been applied to capture key patterns in FC data. In our brain representation learning, we enhance brain network representation by leveraging large language models (LLMs) to extract rich textual descriptions of brain regions, encoding their functional and structural characteristics. These embeddings provide additional semantic information, enriching the graph representations in brain network analysis.
In population networks, neurological disorder prediction is treated as a node classification task. Node features are derived from the upper triangle of FC matrices (Shao et al., 2023) or extracted using methods like recursive feature elimination (RFE) (Yang et al., 2023). Alternatively, GNNs are employed to learn graph representations of FC for node feature embedding (Jiang et al., 2020; Zhang et al., 2023a). Non-imaging data, such as phenotypic features, are integrated by mapping them to edge weights (Huang and Chung, 2020; Zhang et al., 2023a; Kazi et al., 2019) or learning modality-shared and modality-specific representations to construct population networks (Zheng et al., 2022; Chen et al., 2024b). While some recent studies have explored jointly modeling brain-level and population-level graphs to capture both within- and between-subject patterns (Pan et al., 2022; Li and Zhang, 2024; Zeng et al., 2025), we use a two-stage approach to help disentangle subject-specific neural features from broader population-level variations.
@@ 3.1 Weighted Graph Attention in Brain Network
Graph Attention Network (GAT) employs a self-attention mechanism to dynamically learn the attention scores between neighboring brain nodes (Ye and Ji, 2021). At layer $l$, the feature of brain node $i$ is denoted as $\bm{h}_{bi}^{(l)}\in\mathbb{R}^{F}$, and is first linearly transformed by a learnable weight matrix $\bm{W}^{(l)}\in\mathbb{R}^{F^{\prime}\times F}$ to obtain: $\bm{z}_{i}^{(l)}=\bm{W}^{(l)}\bm{h}_{bi}^{(l)}$, where $\bm{W}^{(l)}\in\mathbb{R}^{F^{\prime}\times F}$ is a learnable weight matrix, and $F^{\prime}$ is the dimensionality of the transformed features. The attention score $e_{ij}$ between nodes $i$ and $j$ is computed as: $e_{ij}=\text{LeakyReLU}\left(\bm{a}^{\top}\left[\bm{z}_{i}^{(l)}\|\bm{z}_{j}^{(l)}\right]\right)$, where $\bm{a}\in\mathbb{R}^{2F^{\prime}}$ is a learnable parameter vector, and $\|$ denotes concatenation. Here, $\bm{z}_{i}^{(l)}\|\bm{z}_{j}^{(l)}\in\mathbb{R}^{2F^{\prime}}$ is the concatenated feature vector of brain nodes $i$ and $j$.
@@ 3.2 Heterogeneous Graph Convolution in Population Network
The population graph is defined as $G_{p}=(\mathcal{V}_{p},\mathcal{E}_{p},\bm{X}_{p},\mathcal{C})$, where $\mathcal{V}_{p}$ represents individual nodes, $\mathcal{E}_{p}$ represents population edges, $\bm{X}_{p}\in\mathbb{R}^{|\mathcal{V}_{p}|\times F_{2}}$ is the node feature matrix with $F_{2}$ denoting the input brain feature dimension and $\mathcal{C}$ is a set of conditions that define the relationships between nodes. A heterogeneous graph network (HGCN) (Schlichtkrull et al., 2018; Zhang et al., 2019) is used to propagate information across different relations. Each edge corresponds to a unique relationship under a specific condition $c\in\mathcal{C}$.
@@ 4.1.1 Semantic-Enhanced Brain Graph
In this work, we leverage LLM to derive semantic embeddings of brain regions, capturing both their functional roles and structural properties. For each brain region, prompts are constructed to elicit descriptive text capturing its core neurobiological attributes. These responses are then tokenized and embedded using the text encoder. An illustration of this process is provided in Figure 2. To quantify the semantic relationships between regions, we compute pairwise cosine similarity between their embeddings: $\text{Sim}(u,v)=\frac{\mathbf{s}_{u}\cdot\mathbf{s}_{v}}{|\mathbf{s}_{u}||\mathbf{s}_{v}|}$, where $\mathbf{s}_{u}$ and $\mathbf{s}_{v}$, where $\mathbf{s}_{u}$ and $\mathbf{s}_{v}$ denote the semantic embeddings of brain regions $u$ and $v$, respectively. Based on these similarity scores, we construct edges in the brain graph, using a thresholded similarity score as the edge weight. This incorporation of semantic similarity enables the graph structure to more faithfully reflect the functional and structural relevance between regions.
The semantic similarities serve as a complementary source of information for FC, providing prior knowledge about potential anatomical and functional associations between regions. By integrating this knowledge-driven prior with data-driven FC features, the model is guided to interpret connectivity patterns more effectively and learns meaningful and generalizable representations beyond FC data, which is particularly beneficial in data-scarce or heterogeneous clinical populations.
@@
To ensure consistency and reproducibility in concept generation, we design a structured prompt format grounded in established neuroscience standards, such as the Terminologia Neuroanatomica and the Human Connectome Project. The prompt schema that guides the initial neuroanatomical descriptions is illustrated in Figure 3.
@@
To further enhance the granularity and generalizability of the extracted features, we introduce a secondary prompt tailored to extract hierarchical, ontology-driven concepts for each brain region. This prompt guides the model to produce fine-grained annotations that align with known structural-functional hierarchies. An overview of this concept extraction process is shown in Figure 4. All generated descriptions were manually verified and cross-checked against standard references to ensure terminological consistency
To further enhance the granularity and generalizability of the extracted features, we introduce a secondary, ontology-driven prompt tailored to elicit hierarchical concepts for each brain region. This prompt guides the model to produce fine-grained annotations aligned with known structural–functional hierarchies, as shown in Figure 4. The generated descriptions are manually verified to ensure terminological consistency.
@@ 4.1.2 Feature-to-Group Mapping
As part of the ANR-GAT model, the node feature matrix $\bm{H}_{b}\in\mathbb{R}^{|\mathcal{V}_{b}|\times F^{\prime}}$ is obtained after the WGAT layer. To reassign nodes into new groups based on their feature representations, we introduce a feature-to-group mapping mechanism that projects these features into a shared group space. The group assignment matrix $\bm{M}\in\mathbb{R}^{|\mathcal{V}_{b}|\times N_{g}}$, where $N_{g}$ is the number of groups, is computed by applying a learnable reassignment weight matrix $\bm{W}_{r}\in\mathbb{R}^{F^{\prime}\times N_{g}}$ followed by a softmax operation: $\bm{M}=\text{Softmax}(\bm{H}_{b}\bm{W}_{r})$, where the softmax is applied row-wise to ensure that the group assignment probabilities for each node sum to one.
For each node $i\in\mathcal{V}$, the group assignment $\text{group}_{i}$ is determined by selecting the group with the highest probability from $\bm{M}$ along the group dimension: $\text{group}_{i}=\arg\max_{j}M_{ij},\quad\forall j\in\{1,2,\dots,N_{g}\}.$ With this hard assignment mechanism, the actual group structure is adaptively learned from the data, and the effective number of groups may be smaller than the predefined upper bound $N_{g}$. As illustrated in Figure 1, this mapping is a key component of the ANR-GAT model, designed for brain graph representation learning. Node features are aggregated within each group, combining information from nodes reassigned to the same group. The edges are also remapped between groups, guided by the original graph connectivity and the reassigned groups. This process simplifies the graph structure while preserving key relationships, enabling the generation of new node representations that are passed to subsequent layers of WGAT for further processing. As an intermediate step bridging the WGAT layers, the feature-to-group mapping facilitates hierarchical feature learning and graph structure simplification.
@@ 4.3 Two-Stage Optimization for B2P-GL Framework
The proposed framework comprises two stages. The first stage is devoted to optimizing the representation of the brain graph, while the second stage aims to enhance the classification of disorders.
@@ 5.1 Datasets
Table 1 summarizes the demographic characteristics of the three datasets. ADHD-200 primarily includes children and adolescents, whereas ABIDE I and REST-meta-MDD cover a broader age range extending into adulthood. ABIDE I and ADHD-200 show a clear male predominance, while REST-meta-MDD has a more balanced gender distribution. All datasets are collected from multiple sites, contributing to demographic and acquisition variability. ABIDE I includes 850 participants from 20 sites, ADHD-200 consists of 938 participants from 7 sites, and the REST-meta-MDD comprises 1,567 participants across 16 sites. While both ABIDE I and ADHD-200 include Full-Scale, Performance, and Verbal IQ scores, a proportion of these data are missing, with 22% missing in ABIDE I and 26.3% in ADHD-200. REST-meta-MDD includes years of education, which may serve as a proxy for cognitive and socioeconomic background.
@@
5.2 Graph Definition in Brain Graph
To construct the brain graph used in our framework, we define the nodes and edges for each dataset as follows. For the ABIDE and ADHD-200 datasets, we used the Harvard-Oxford (HO) atlas (Desikan et al., 2006). Although it defines 111 ROIs, we excluded one unlabeled region, resulting in 110 ROIs. For the MDD datasets, we used the AAL atlas with its 116 ROIs(Tzourio-Mazoyer et al., 2002). Features were calculated as the Pearson correlation coefficients between every pair of ROIs within each atlas.
For the text encoder, we used BioBERT (Lee et al., 2019), a BERT-based language model pre-trained on biomedical literature. For each region, embeddings were generated by averaging the token embeddings from the last four hidden layers of the model that effectively capture nuanced contextual information. Edges between nodes were subsequently defined based on the similarity of these region embeddings.
5.3 Processing of Phenotypic Data
To enable meaningful categorical comparisons, age was stratified into six groups to ensure balanced sample sizes. Phenotypic variables, including site, gender, and age group, were encoded using one-hot encoding. Given the critical role of intelligence quotient (IQ) in neurocognitive function (Nisbett et al., 2012), particularly in neurodevelopmental disorders (Wolff et al., 2022), we incorporated IQ data into the phenotypic features, accounting for variations in test types. IQ metrics considered included Full Scale IQ, Verbal IQ, and Performance IQ. The type of IQ test administered was also included to address potential variations in IQ distributions. For missing IQ data, we imputed the mean value for each test type to ensure consistency and maintain the accuracy of the dataset.
@@ 5.4 Experimental Settings
Models were implemented and evaluated within a PyTorch environment using an NVIDIA A800 GPU with 80 GB memory. Table 2 summarizes the hyperparameter configurations for the two stages of the B2P-GL model across the ABIDE I, ADHD-200, and REST-meta-MDD datasets. The first stage adopts the ANR-GAT architecture, while the second stage is based on the HGCN framework. The table outlines dataset-specific settings for each stage, including the optimizer, learning rate, batch size, number of training epochs, hidden dimensions, and selected phenotypic variables.
@@
learning rate $0\times 10^{-4} 1.00×1$$0\times 10^{-4}$ $1.00\times 10^{-4}$
@@
learning rate $0\times 10^{-2} 1.00×1$$0\times 10^{-2}$ $1.00\times 10^{-2}$
@@
5.5 Model Performance and Representation Analysis
@@
Figure 5: Visualization of brain feature embeddings using t-SNE. Compared with stage 1 embeddings derived from individual graphs and stage 2 embeddings updated through HGCN without phenotypic fusion.
To provide a more direct demonstration of the effectiveness of condition-based graph construction and HGCN in stage 2, we further visualized the feature embeddings using t-SNE. We compared the brain feature embeddings learned from stage 1 with those updated through HGCN without phenotypic fusion. In Figure 5, the class distance increased markedly from 0.6769 in stage 1 to 16.2326 in stage 2, indicating that the embeddings obtained in the population graph capture more disease-relevant information beyond site and demographic effects.
@@ 5.6.1 Effectiveness of Semantic Similarity in Edge
In addition, we further evaluated two strategies for computing semantic similarity: (1) lexical embeddings of region names and (2) contextualized embeddings incorporating their concepts. Subsequently, we compared BERT-base (Devlin et al., 2018), a pre-trained language model designed for general tasks, and BioBERT, a biomedical domain-adapted variant.
As shown in Table 5, incorporating contextualized concepts led to consistent performance gains. For BERT-base, concept-based embeddings improved accuracy by 0.6%. BioBERT also benefited from concept augmentation, with a 0.4% increase. When comparing models under the same input setting (i.e., concepts), BioBERT outperformed BERT-base by 0.3% in accuracy and 0.4% in AUC, indicating its advantage as a more suitable choice for neurological applications. These results highlight the importance of both domain-adapted language models and concept-level augmentation for capturing meaningful semantic similarity between brain regions.
@@
Names BERT 72.8 $\pm$ 0.6 72.3 $\pm$ 0.6 80.2 $\pm$ 1.8 64.4 $\pm$ 2.4
BioBERT 73.3 $\pm$ 0.8 72.9 $\pm$ 0.8 79.2 $\pm$ 1.5 66.5 $\pm$ 1.6
Concepts BERT 73.4 $\pm$ 0.5 72.8 $\pm$ 0.6 81.7 $\pm$ 2.0 64.0 $\pm$ 2.5
BioBERT 73.7 $\pm$ 0.6 73.2 $\pm$ 0.6 80.5 $\pm$ 1.8 65.9 $\pm$ 2.1